# Compton Form Factor Extraction using Quantum Deep Neural Networks

### v2

@@ A.3 Tables
Table 19: Results from the optimized CFF extraction from the real experimental data subset performed in Sec. VII. For each set of kinematics, the mean and standard deviation of the FQDNN-predicted CFFs are listed.
Set $k$ $Q^{2}$ $x_{B}$ $t$ Mean $\Re e\mathcal{H}$ Std $\Re e\mathcal{H}$ Mean $\Re e\mathcal{E}$ Std $\Re e\mathcal{E}$ Mean $\Re e\widetilde{\mathcal{H}}$ Std $\Re e\widetilde{\mathcal{H}}$ Mean $DVCS$ Std $DVCS$
108 5.750 1.790 0.244 -0.110 -1.304 0.714 -1.604 0.798 -1.548 0.848 -0.0137 0.0200
109 5.750 1.790 0.244 -0.150 -0.222 0.498 -0.540 0.825 0.058 0.757 0.0201 0.0113
110 5.750 1.790 0.244 -0.200 -0.934 0.503 -0.962 0.544 -1.448 0.726 0.0080 0.0085
111 5.750 1.790 0.244 -0.260 -0.577 0.543 -0.724 0.635 -0.208 1.085 0.0196 0.0076
112 5.750 1.800 0.244 -0.340 -0.465 0.905 -0.731 0.912 0.160 0.836 0.0222 0.0080
113 5.750 1.800 0.244 -0.450 -0.508 1.640 -0.400 1.404 -0.240 0.596 0.0166 0.0090
114 5.750 2.120 0.244 -0.110 1.270 0.826 1.376 0.855 1.405 0.901 -0.0234 0.0170
115 5.750 2.120 0.245 -0.150 2.950 0.990 2.909 0.908 2.937 1.127 0.0406 0.0136
116 5.750 2.120 0.245 -0.200 4.590 1.532 3.880 1.371 3.496 1.437 0.0453 0.0176
117 5.750 2.120 0.245 -0.260 5.603 1.775 4.825 1.587 3.557 1.468 0.0573 0.0170
118 5.750 2.120 0.245 -0.340 9.494 3.111 8.464 2.769 5.397 2.212 0.0723 0.0223
119 5.750 2.120 0.245 -0.450 0.538 1.557 0.588 1.438 0.119 0.931 0.0193 0.0101
120 5.750 1.940 0.274 -0.110 -0.594 1.161 -0.973 1.988 -1.083 1.833 0.0261 0.0280
121 5.750 1.940 0.274 -0.150 -1.234 0.626 -1.707 0.754 -1.894 0.795 0.0022 0.0119
122 5.750 1.940 0.274 -0.200 -0.700 0.439 -1.177 0.575 -0.523 0.674 0.0028 0.0078
123 5.750 1.940 0.274 -0.260 -1.132 0.460 -1.286 0.491 -1.729 0.570 0.0091 0.0069
124 5.750 1.940 0.275 -0.340 -0.152 0.504 -0.218 0.573 0.392 1.135 0.0086 0.0052
125 5.750 1.940 0.275 -0.450 -0.159 0.756 -0.129 0.919 -0.305 0.613 0.0133 0.0040
126 5.750 2.350 0.274 -0.110 0.911 1.706 1.269 2.371 1.203 2.245 -0.0420 0.0366
127 5.750 2.350 0.275 -0.150 0.580 0.911 0.701 1.094 0.718 1.204 0.0038 0.0105
128 5.750 2.350 0.275 -0.200 1.874 0.899 1.975 0.874 2.159 1.023 0.0150 0.0073
129 5.750 2.350 0.275 -0.260 3.161 1.189 3.095 1.138 2.886 1.323 0.0288 0.0077
130 5.750 2.350 0.275 -0.340 3.104 1.320 2.424 1.122 1.522 0.858 0.0246 0.0067
131 5.750 2.350 0.275 -0.450 7.237 2.637 6.298 2.351 3.682 1.580 0.0340 0.0102
133 5.750 2.100 0.304 -0.150 -2.007 1.511 -3.181 2.303 -3.098 2.231 0.0263 0.0319
134 5.750 2.100 0.304 -0.200 -0.690 0.453 -1.249 0.698 -1.231 0.619 0.0040 0.0062
135 5.750 2.100 0.304 -0.260 -0.651 0.437 -0.938 0.567 -1.174 0.719 0.0058 0.0059
136 5.750 2.100 0.304 -0.340 -0.662 0.322 -0.802 0.368 -1.464 0.499 0.0053 0.0040
137 5.750 2.100 0.305 -0.450 -0.329 0.390 -0.368 0.424 -0.547 0.845 0.0116 0.0029
139 5.750 2.580 0.304 -0.150 -0.852 2.441 -1.130 3.231 -1.087 3.148 0.0128 0.0421
140 5.750 2.580 0.304 -0.200 0.094 0.545 0.143 0.847 0.085 0.918 -0.0003 0.0049
141 5.750 2.580 0.305 -0.260 2.103 0.771 2.269 0.740 2.566 0.799 0.0100 0.0039
142 5.750 2.580 0.305 -0.340 1.524 0.932 1.620 0.909 1.085 0.907 0.0106 0.0038
143 5.750 2.580 0.305 -0.450 2.056 1.553 1.510 1.189 0.790 0.741 0.0135 0.0042
145 5.750 2.230 0.335 -0.200 -0.086 0.381 -0.197 0.933 -0.271 1.036 -0.0060 0.0048
146 5.750 2.230 0.335 -0.260 -0.658 0.456 -1.031 0.669 -1.619 0.773 0.0080 0.0059
147 5.750 2.230 0.335 -0.340 -0.429 0.393 -0.548 0.484 -0.867 0.780 0.0086 0.0040
148 5.750 2.230 0.335 -0.450 -1.064 0.311 -1.210 0.338 -2.108 0.395 0.0044 0.0026
150 5.750 2.780 0.335 -0.200 -0.914 0.635 -1.603 0.866 -1.505 0.888 0.0026 0.0054
151 5.750 2.780 0.335 -0.260 -0.052 0.385 -0.081 0.768 -0.223 0.864 0.0015 0.0027
152 5.750 2.780 0.335 -0.340 1.135 0.615 1.442 0.658 1.700 0.760 0.0075 0.0022
153 5.750 2.780 0.335 -0.450 0.784 0.845 0.989 0.995 0.429 0.650 0.0060 0.0020
154 5.750 2.350 0.365 -0.200 -1.929 1.139 -5.204 2.382 -4.058 2.066 0.0097 0.0110
155 5.750 2.350 0.365 -0.260 -0.902 0.543 -1.813 0.936 -1.961 0.998 0.0047 0.0057
156 5.750 2.350 0.365 -0.340 -0.340 0.412 -0.507 0.591 -0.836 0.943 0.0061 0.0040
157 5.750 2.350 0.365 -0.450 -0.373 0.284 -0.462 0.343 -0.983 0.606 0.0036 0.0027
158 5.750 2.970 0.364 -0.200 -0.330 0.875 -0.857 2.353 -0.649 1.728 -0.0014 0.0055
159 5.750 2.970 0.365 -0.260 0.007 0.331 -0.015 0.970 0.083 1.067 0.0037 0.0026
160 5.750 2.970 0.365 -0.340 0.583 0.400 1.137 0.603 1.293 0.707 0.0021 0.0017
161 5.750 2.970 0.365 -0.450 0.976 0.391 1.283 0.444 1.943 0.444 0.0036 0.0011
163 5.750 2.480 0.400 -0.260 0.140 0.521 0.375 1.376 0.293 1.123 -0.0001 0.0047
164 5.750 2.480 0.399 -0.340 -0.193 0.524 -0.323 0.821 -0.434 1.061 0.0057 0.0045
165 5.750 2.480 0.399 -0.450 -0.699 0.312 -0.905 0.385 -1.481 0.574 0.0018 0.0024
166 5.750 3.120 0.391 -0.210 -1.311 1.796 -3.921 4.696 -2.405 3.127 0.0191 0.0134
167 5.750 3.180 0.399 -0.260 -1.064 0.742 -2.757 1.586 -2.147 1.324 0.0002 0.0038
168 5.750 3.180 0.400 -0.340 -0.347 0.273 -1.059 0.645 -1.434 0.731 0.0023 0.0015
169 5.750 3.180 0.400 -0.450 0.073 0.139 0.155 0.269 0.189 0.795 0.0032 0.0009
171 5.750 2.960 0.449 -0.340 0.248 0.424 0.757 1.164 0.549 0.867 0.0010 0.0027
172 5.750 3.050 0.461 -0.450 -0.114 0.186 -0.276 0.400 -0.411 0.521 0.0016 0.0009
173 5.750 3.500 0.430 -0.280 -1.382 1.844 -3.202 3.750 -2.327 2.896 -0.0006 0.0098
174 5.750 3.630 0.451 -0.340 -0.072 0.345 -0.256 1.167 -0.168 0.723 0.0006 0.0016
175 5.750 3.770 0.475 -0.450 0.108 0.174 0.279 0.440 0.275 0.405 0.0010 0.0006

### v3

@@ V.1.2 CFF Extraction Test
V.2 Fully Optimized QDNN
V.2.1 Full QDNN Model
To systematically explore quantum-specific optimizations, we explore three new QDNN models (Models 1-3), each incorporating targeted modifications designed to enhance specific aspects of the architecture with the goal of improving stability and reducing errors. Our baseline model, the Basic QDNN—introduced earlier in Sec. V.1—is a straightforward quantum adaptation of the CDNN. Although not fully optimized, it serves as a proof of concept for achieving a quantum advantage in CFF extraction and provides a valuable foundation for identifying potential improvements. The QDNN algorithmic error improvements observed from Models 1-3 are more comparable to the algorithmic errors from the CDNN extraction (see Fig. 6) and suggest key strategies for reducing algorithmic error and improving precision in quantum deep learning models for CFF extraction. Specifically, by refining the entanglement structure, implementing layer-wise training, and adjusting initialization schemes, we demonstrate that targeted modifications can significantly enhance training stability and overall model reliability. Therefore, to construct our FQDNN, we incorporate all the improvements from Models 1-3, including tunable entanglement range (Model 1), layer-wise training (Model 2), and depth-scaled, larger-angle initialization (Models 2 and 3), while maintaining the baseline architecture from the Basic QDNN, including a width of 6 qubits and classical pre- and post-processing layers.
To test the FQDNN, we perform the same error analysis used on Models 1-3 and display the results in Table 4. Incorporating features from Models 1-3 allows the FQDNN to perform better overall than each of the individual models, with improved algorithmic error and precision over all 4 CFFs. With these example kinematics, the FQDNN also exhibits improved methodological error for $\Re e\mathcal{E}$, $\Re e\widetilde{\mathcal{H}}$, and $DVCS$ and improved accuracy for $\Re e\widetilde{\mathcal{H}}$ and $DVCS$.
Model 1 modifies the strongly entangling layers by adding an entanglement range parameter, limiting interactions to nearest neighbors in some cases rather than fully entangling all qubits at every layer. This aims to balance expressivity and stability, reducing excessive parameter correlations that can destabilize optimization. Model 2 focuses on addressing the barren plateau problem, a QDNN-unique issue characterized by gradients that vanish exponentially as the circuit depth increases. To mitigate this, Model 2 employs layerwise training, where the network begins with a shallow quantum circuit and progressively increases the number of layers over training. Additionally, it introduces depth-scaled initialization, ensuring that each layer’s parameter magnitudes decrease as depth increases. This prevents excessive parameter magnitudes in deeper layers while maintaining trainable gradients throughout optimization. Model 3 refines the small-angle initialization used in the Basic QDNN, which is already known to help prevent barren plateaus at initialization but may overly restrict the search space, leading to increased error. Instead of reducing the initial weight magnitudes across layers, Model 3 rescales small-angle initialization to maintain better trainability without completely restricting parameter exploration.
To test these new models, we perform the CFF extraction on 1000 noisy replicas of Set 144 (see Fig. 3) using each of these four models and display the algorithmic error results in Table 4. Comparing the results of the Basic QDNN with Fig. 6, we observe that the errors for this kinematic bin align relatively well with the average errors of the extracted CFFs across all bins. Performing the extraction with the improved QDNN models, the algorithmic error and precision improve significantly for all three models. Table 4 indicates that Model 1 produces the lowest algorithmic errors for $\Re e\mathcal{H}$, $\Re e\mathcal{E}$, and $\Re e\widetilde{\mathcal{H}}$, while Model 2 produces the lowest algorithmic error for $DVCS$. The improved precision follows a similar pattern. In this example, we find that these new models significantly improve accuracy and methodological error for $\Re e\widetilde{\mathcal{H}}$, but there is no overall trend for these two metrics.
(a) CDNN
(b) Basic QDNN
(c) Full QDNN
Figure 8: Predicted cross sections from the (a) CDNN, (b) Basic QDNN, and (c) Full QDNN fits of noisy replicas of cross section pseudodata (red points) generated from kinematics $k=8.847$ GeV, $Q^{2}=5.36$ GeV2, $x_{B}=0.485$, $t=0.50851$ GeV2 (Set 48). In this example, the FQDNN exhibits a dramatic improvement from the CDNN and BQDNN in terms of proximity to the true cross section and a small improvement in accuracy.
V.2.2 CFF Extraction Results
These results demonstrate that QDNNs provide a powerful framework for extracting CFFs from DVCS data and, when the quantum architecture is fully optimized, generally achieve smaller uncertainties than CDNNs. The results shown in Figs. 6 and 7 represent means evaluated over the full kinematic range of the data. While the FQDNN typically yields CFFs that better reproduce the true cross section, the CDNN can outperform the QDNN in certain kinematic regions or error regimes. This naturally raises the question of how to select the optimal algorithm for a given dataset, kinematic configuration, or experimental uncertainty. To address this, we introduce a novel quantum qualifier metric that predicts a priori whether a CDNN or FQDNN is better suited for CFF extraction under specified experimental conditions by identifying regimes in which QDNNs are expected to outperform CDNNs.
(a) $0\sigma_{F}$ experimental error
(b) $2\sigma_{F}$ experimental error
Figure 9: Predicted cross section from the CDNN and QDNN fits of cross section pseudodata (red points) generated from kinematics $k=5.75$ GeV, $Q^{2}=2.012$ GeV2, $x_{B}=0.378$, t=0.192t=-0.192 GeV2 and CFFs $\Re e\mathcal{H}\approx-0.808$, $\Re e\mathcal{E}\approx-1.229$, $\Re e\widetilde{\mathcal{H}}\approx-1.360$, and $DVCS=0.0114$. This example comparison illustrates the general trend of increased quantum outperformance $\Xi_{DVCS}$ with greater experimental error.
VI DVCS Quantum Qualifier
In this section, we construct a quantum qualifier $\hat{\Xi}_{DVCS}$ tailored specifically for DVCS data. The goal of the quantum qualifier is to determine whether the CDNN or FQDNN (described in Sec. V.2) should be used for the extraction of CFFs from a given set of experimental DVCS data, based solely on its functional characteristics and error. The quantum qualifier is designed based on CDNN and FQDNN extractions from systematically constructed pseudodata with physically realistic kinematics, CFFs, and error. Then, the quantum qualifier is tested, and in the following section (Sec. VII), it is used to perform a preliminary check to determine which algorithm (QDNN or CDNN) is optimal for that particular kinematic set for maximal extraction of CFF information.
VI.1 Pseudodata Generation
To construct the DVCS quantum qualifier, we generate a controlled ensemble of pseudodata based on the kinematic values of the experimental measurements. As a first step, we perform a baseline CDNN extraction on 1000 replicas of each experimental data set, sampling each cross-section point within its quoted uncertainty. This provides a physically grounded range of CFF values—typically between 5-5 and 55—that serves as the foundation for the pseudodata construction.
Using these CDNN-extracted CFFs, we systematically generate pseudodata for each of the 195 kinematic sets. For every CFF, we select five values evenly spaced within the 95% confidence interval obtained from the baseline extraction, and for each cross-section point we impose one of four error scalings: 0, $0.5\sigma_{F}$, $1\sigma_{F}$, or $2\sigma_{F}$, where $\sigma_{F}$ is the experimental uncertainty. This results in $5^{4}\cdot 4=2500$ pseudodata replicas per kinematic point, providing a comprehensive and physically consistent dataset for training the DVCS quantum qualifier.
VI.2 Quantum Qualifier
The DVCS quantum qualifier $\hat{\Xi}_{DVCS}$ is designed to estimate the underlying quantum outperformance metric $\Xi_{DVCS}$. Defining this metric requires the quantity $M_{DVCS}$, given by:
$M_{DVCS}=\int_{\phi_{\text{min}}}^{\phi_{\text{max}}}|F_{\text{DNN}}(\phi)-F_{\text{true}}(\phi)|d\phi,$ (30)
$\Xi_{DVCS}=M_{DVCS}^{\text{CDNN}}/M_{DVCS}^{\text{QDNN}}-1,$ (31)
[…]
where
$\bar{y}=\frac{1}{n}\sum_{i=1}^{n}y_{i}.$ (33)
[…]
with $N$ the total number of $(i,r)$ pairs. Varying $\overline{\epsilon}_{s}$ across the pseudodata sets from 0 to 0.6 allows us to study the impact of increasing experimental uncertainty. We find a strong positive correlation between $\Xi_{DVCS}$ and $\overline{\epsilon}_{s}$, well-approximated by the linear relation $\Xi_{DVCS}=2.00\overline{\epsilon}_{s}-0.66$.
This trend is illustrated in Fig. 9. When there is no experimental error, the CDNN achieves a nearly perfect fit (Fig. 9a), whereas the FQDNN underperforms slightly. However, under high experimental uncertainty (Fig. 9b), the CDNN performance degrades significantly, while the FQDNN remains robust—highlighting the increased quantum outperformance with increasing error.
[…]
A positive $\hat{\Xi}_{DVCS}$ suggests the QDNN will likely outperform, with its magnitude indicating the strength of correlation with out-performance. While $\hat{\Xi}_{DVCS}$ guides predictions of QDNN advantage, it does not precisely determine $\Xi_{DVCS}$, which depends on data structure and kinematics. A linear fit of $\Xi_{DVCS}=1.0004\hat{\Xi}_{DVCS}-0.0042$ ($R^{2}=0.42$), based on average $\hat{\Xi}_{DVCS}$ across kinematic sets and noise level $s$ (excluding extreme outliers), confirms approximate agreement with quantum out-performance.
[…]
This simple utility provides the means of checking experimental data sets to determine which algorithm type would perform best and can be used to determine the optimal fit strategy for any given set of experimental data. Note that $\hat{\Xi}_{DVCS}$ is a selection heuristic, not a performance guaranty. For our pseudodata set the conclusion made by using this heuristic fails on less than 3% of the test data. With the quantum qualifier defined and its predictive power established, we apply our optimized CFF extraction method to real experimental data for improved extraction.
VII Optimized CFF Extraction from Experimental Data
This section demonstrates the practical application of the DVCS quantum qualifier $\hat{\Xi}_{DVCS}$ for optimized CFF extraction on a subset of real experimental data (see Sec. IV), validating our QDNN approach in a realistic setting beyond pseudodata. We select a subset where $\hat{\Xi}_{DVCS}$ predicts quantum outperformance in most kinematics to evaluate the quantum advantage. This subset has enough data for a good demonstration, but with a variety of experimental errors.
CFFs $a$ $b$ $c$ $d$ $e$ $f$
$\Re e\mathcal{H}$ -8.13 1.82 35.26 25.37 6.26 3.20
$\Re e\mathcal{E}$ 6.92 -5.64 0.81 0.98 4.03 49.71
$\Re e\widetilde{\mathcal{H}} -$8.51 1.72 31.11 22.49 6.09 4.77
DVCSDVCS 0.45 -0.45 4.40 2.91 0.13 0.08
(a) $\Re e\mathcal{H}$ vs. t-t with $x_{B}=0.275$
(b) $\Re e\mathcal{E}$ vs. t-t with $x_{B}=0.365$
(c) $\Re e\widetilde{\mathcal{H}}$ vs. t-t with $x_{B}=0.399$
(d) DVCSDVCS vs. t-t with $x_{B}=0.244$
Figure 10: Example comparisons between the initial CDNN-extracted CFFs (blue with error bars) and the CFFs produced from the generating function [Eq. (19)] using the fitted parameters in Table 5 (red). The CFFs are plotted against t-t for select values of $x_{B}$ in the subset of all experimental data specified in the text.
VII.1 Realistic Pseudodata Generation
We choose to examine a specific subset of the experimental data (Sets 108 to 175) that has kinematics $k=5.75$ GeV, $Q^{2}$ from 1.79 GeV2 to 3.77 GeV2, $x_{B}$ from $0.244$ to $0.475$, and $t$ from 0.11-0.11 GeV2 to 0.45-0.45 GeV2. This subset selects 1145 points in a physically relevant region of phase space covered by JLab experiments (Sec. IV), where the DVCS cross-section data has a range of experimental errors.
Using the improved generating function as a stand-in for the “true” kinematic dependence of the CFFs, we construct a new set of realistic pseudodata. Following the procedure in Sec. IV, we use the fitted function to generate the CFFs for each experimental kinematic setting. These CFFs are then used to compute the DVCS cross section at each $\phi$ point. To simulate realistic measurement uncertainties, we add Gaussian noise to each cross-section value based on the established experimental error. Finally, for each kinematic point, we generate 1000 replicas of noisy cross-section data by resampling within the experimental uncertainty. These realistic pseudodata sets are then used for the tuning and evaluation of the DNNs.
(a) $\Re e\mathcal{H}$ vs t-t with $x_{B}=0.244$
(b) $\Re e\mathcal{E}$ vs t-t with $x_{B}=0.304$
(c) $\Re e\widetilde{\mathcal{H}}$ vs t-t with $x_{B}=0.399$
(d) DVCSDVCS vs t-t with $x_{B}=0.335$
(a) $\Re e\mathcal{H}$ vs. $x_{B}$ and t-t
(b) $\Re e\mathcal{E}$ vs. $x_{B}$ and t-t
(c) $\Re e\widetilde{\mathcal{H}}$ vs. $x_{B}$ and t-t
Figure 12: Three-dimensional surface plots of (a) $\Re e\mathcal{H}$, (b) $\Re e\mathcal{E}$, (c) $\Re e\widetilde{\mathcal{H}}$ vs. $x_{B}$ and t-t, generated showing the mean of the global fit result for three values of $Q^{2}$ 1.0, 4.0, 8.0 GeV2 with 1.0 GeV2 on the bottom in each case.
VII.2 Realistic Pseudodata Extraction
The quantum qualifier DVCS $\hat{\Xi}_{DVCS}$ [Eq. (35)] indicates that the FQDNN should outperform nearly all ($96\%$) of the pseudodata sets. This is confirmed by the realistic pseudodata extraction, which shows overwhelming quantum outperformance. In Fig. 11, we plot example comparisons of the extracted CFFs by the tuned CDNN and FQDNN. As predicted by the DVCS quantum qualifier, the FQDNN demonstrates improved accuracy and precision across the entire range of kinematic values covered by the realistic pseudodata sets, effectively capturing the trend in the CFFs.
(a) $\Re e\mathcal{H}$ vs t-t with $x_{B}=0.365$
(b) $\Re e\mathcal{E}$ vs t-t with $x_{B}=0.275$
(c) $\Re e\widetilde{\mathcal{H}}$ vs t-t with $x_{B}=0.305$
Figure 13: Example comparisons between the FQDNN-extracted CFFs from the real experimental data and the KM15 prediction. The FQDNN mean is show as the blue line with shaded $1\sigma$ error band, and the KM15 curve is shown in red.
VII.3 Experimental Data Extraction
The construction of $\hat{\Xi}_{\text{DVCS}}$ allows us to go back and evaluate the entire experimental dataset in Table 1 and fit each local dataset using either the FQDNN or the standard CDNN for best results. We find that over 60% of the data benefits from the FQDNN approach. Roughly 20% of the data fits equally well with either algorithm. The remaining 20% fits better with the CDNN approach. All of the optimally fitted CFFs (either with FQDNN or CDNN) are subsequently used in a standard DNN-based global fit to obtain a continuous CFF model throughout the measured phase space.
[…]
Experimental and algorithmic uncertainties are propagated simultaneously by training 1000 independent DNNs on bootstrap replicas of the data set. At any kinematic point $(x_{B},t,Q^{2})$, the central CFF value is taken as the ensemble mean $\overline{\mathcal{F}}_{\text{DNN}}$, and the total uncertainty is given by the standard deviation of the replica predictions,
$\sigma(x_{B},t,Q^{2})=\sqrt{\frac{1}{N-1}\sum_{j=1}^{N}\left(\mathcal{F}^{j}_{\text{DNN}}-\overline{\mathcal{F}}_{\text{DNN}}\right)^{2}},$ (36)
with $N=1000$. The resulting continuous CFF surfaces are shown as functions of $x_{B}$ and t-t in Fig. 12.
For consistency checks, we compare our resulting global fit to the KM15 model [81], a modern global fit for DVCS that parametrizes the dominant CFFs using a flexible GPD ansatz. It builds on earlier KM10b [80] and KMM12b [79] fits and incorporates updated DVCS cross section and asymmetry data, including the earlier, but incomplete data of same measurements from Hall A E07-007 and E00-110 [31] and CLAS Hall B e1-DVCS1 [52, 50] as in Table 1. The model includes both $H$ and $E$ GPDs, with a simplified functional form that allows for control over $t$-dependence and skewness effects. It assumes factorized $t$-dependence and parametrizes the GPDs using Regge-inspired profiles and a spectral representation consistent with polynomiality constraints. The KM15 fit reveals that the skewness effect in GPD $H$ is small, and the $t$-dependence of $E$ is flatter than that of $H$, aligning with expectations from dispersion relations. The model is constrained by experimental observables through Fourier harmonics of the cross section, with sensitivity enhanced by weighting techniques. Fig. 13 shows that the our global replica mean as the blue line is consistent within $1\sigma$ error bands with the KM15 model (red line) for $\Re e\mathcal{E}$, and $\Re e\widetilde{\mathcal{H}}$, but deviates with a clear increasing trend in t-t for $\Re e\mathcal{H}$.
VII.4 Comparison to Previous Extractions
[…]
Figure 14: Comparison of our global DNN predictions with previous extractions from experimental data.
[…]
The improvement introduced by the QDNN manifests itself not merely as a narrower uncertainty band but as a qualitatively different behavior of the uncertainty itself. Regions of limited experimental coverage show expected widening, yet elsewhere the uncertainties shrink markedly—reflecting the ability of QDNN to exploit latent correlations to stabilize the learning of $\Re e\,\mathcal{H}$, $\Re e\,\mathcal{E}$, and $\Re e\,\widetilde{\mathcal{H}}$. This intrinsic stabilization emerges without imposing external physical constraints, indicating that the quantum-encoded latent space acts as a self-regularizing prior.
From a methodological standpoint, the QDNN therefore extends the concept of a physics-informed network beyond classical architectures: it learns the complex-valued structure of the DVCS amplitude directly, rather than inferring it indirectly from nonlinear regressions on the observables. The result is a more stable convergence behavior, improved generalization to unseen kinematics, and a reduction in model bias that persists even when trained on the least-constraining dataset.
Overall, the QDNN approach demonstrates that quantum-differentiable representations can reproduce—and in several cases exceed—the precision of prior DNN global fits that relied on a richer set of observables. This establishes the QDNN as a robust and scalable framework for future global extractions, capable of incorporating additional polarization observables without sacrificing interpretability or stability.
VIII Conclusion
Within a common BKM10 twist-2 forward model, we have presented the first extraction of Compton form factors (CFFs) from experimental DVCS cross sections using a quantum deep neural network (QDNN), together with a controlled, like-for-like benchmark against a classical deep neural network (CDNN) for this application. Across extensive closure tests on kinematically realistic pseudodata and validation on experimental subsets, optimized QDNNs achieve tighter CFF uncertainties and, in many regimes, comparable or improved accuracy relative to CDNNs. The gains are most pronounced in sparse or high-uncertainty kinematic regions and for more complex (more nonlinear) $\phi$-dependence of the measured cross section.
Methodologically, we identified and mitigated optimization instabilities that are specific to quantum-circuit models. In particular, tunable entanglement range, layer-wise depth growth, and depth-scaled initialization stabilize training and reduce algorithmic variability, bringing algorithmic errors to a level comparable to classical baselines while producing visibly tighter uncertainty bands in the reconstructed cross sections.
A practical outcome of this study is the DVCS quantum qualifier $\hat{\Xi}_{\mathrm{DVCS}}$: a simple, data-driven selection rule built from measurable features of a dataset (the average scaled experimental error and a nonlinearity metric) that predicts a priori when a QDNN is expected to outperform a CDNN. In our tests, the qualifier successfully guided per-bin model selection and enabled an optimized strategy in which the best-performing local extraction (QDNN or CDNN) is retained for each kinematic bin.
Applying this qualifier-guided approach to the full set of experimental data yields a set of high-quality local CFF extractions that we then use to train a global DNN parameterization over the measured phase space. The resulting trends are consistent with established global analyses for $\Re e\,\mathcal{E}$ and $\Re e\,\widetilde{\mathcal{H}}$ within uncertainties, while $\Re e\,\mathcal{H}$ exhibits a systematic difference in its $t$-dependence. Overall, the combined local and global strategy delivers consistently reduced uncertainties across the kinematic range compared to representative previous extractions.
Looking ahead, the framework is naturally extensible to full eight-CFF and polarization-dependent extractions, and to increased theory fidelity through QCD evolution and next-to-leading-order effects. Although the present results are obtained with simulator-based quantum circuits, the architectures are hardware-compatible, providing a direct path to native execution as quantum processors mature. The framework and software used in this analysis are available at the UVA Spin Physics GitHub [57].
Acknowledgments